OpenStudy (anonymous):

Prove the identity: (sin^2theta) x (csc^2theta) = sin^2theta +cos^2theta

OpenStudy (freckles):

First of all what is sin^2(x)+cos^2(x)=?

OpenStudy (freckles):

This is the most basic trig identity also know as a Pythagoras identity

OpenStudy (freckles):

And actually this identity has some restrictions

OpenStudy (homeworksucks):

\[\sin^{2} \theta \csc ^{2} \theta=\sin^{2} \theta +\cos^{2} \theta\]

OpenStudy (anonymous):

I honestly have no idea how to prove identities.. I;m sorry @freckles & @Homeworksucks

OpenStudy (freckles):

You should know sin^2(x)+cos^2(x)=1 and csc(x) =1/sin(x) Use these here

OpenStudy (anonymous):

I don't really know anything, I'm new to this trig thing and my teacher is not helping at all. the way he explains it is too complicated

OpenStudy (aravindg):

You can't get to an answer unless you make a good attempt. csc theta=1/sin theta Substitute that here and tell me what you get (sin^2theta) x (csc^2theta)